OpenStudy (anonymous):

If f(x) = x2, which of the following describes the graph of f(x) - 5?

OpenStudy (anonymous):

please HELP !

OpenStudy (ahsome):

Equation: \[f(x)=x^2\]

OpenStudy (anonymous):

yeah but how to graph of f(x)-5 ? The graph of f(x) - 5 is a vertical shift of f(x) = x2 five units down. The graph of f(x) - 5 is a vertical shift of f(x) = x2 five units up. The graph of f(x) - 5 is a horizontal shift of f(x) = x2 five units to the left. The graph of f(x) - 5 is a horizontal shift of f(x) = x2 five units to the right.

OpenStudy (anonymous):

@Ahsome

OpenStudy (mathmath333):

\(\large\tt \color{black}{plot~~the~~graph~~y=x^2-5 }\) \(\large\tt \color{black}{plot~~the~~points~~\\(1,-4)(-1,4)(-2,-1)(2,-1)(-3,4)(3,4) and (0,5) }\)